You are a caterer and you need to make a 22.5 pound fruit salad. The ratio is 2 parts pineapple, 4 parts strawberries and 6 parts grapes. Using your knowledge of ratios and proportions how much of each type of fruit is in the salad?
Found 2 solutions by stanbon, g6qwerty:
Answer by stanbon(75887)

You are a caterer and you need to make a 22.5 pound fruit salad. The ratio is 2 parts pineapple, 4 parts strawberries and 6 parts grapes. Using your knowledge of ratios and proportions how much of each type of fruit is in the salad?
------
2:4 is the same as 2x:4x where 2x is pineapple and 4x is strawberries.
----
Equation:
2x + 4x = 22.5
6x = 22.5
x = 3.75
-----------------
pineapple = 2x = 7.5 lbs
strawberries = 4x = 15 lbs
================================
Cheers,
Stan H.
==================

Answer by g6qwerty(1)

2:4:6 = 1 complete fruit salad
so 2+4+6=16
2/16*22.5=3.75lb pineapple
4/16*22.5=7.5lb strawberries
6/12*22.5=11.25lb grapes
Previous answer missed the grapes in the above question.